OpenStudy (worldwarkatyy):

Help!!! Will give medal!!! 1. Solve the following equation and check the solution. 4/x-2 = x-1/x-2 2. Determine if the hyperbola is vertical or horizontal. Then find the vertices, foci, and asymptotes. Y^2/25 = x^2/100 =1

OpenStudy (mertsj):

Is this the problem: \[\frac{4}{x-2}=\frac{x-1}{x-2}\]

OpenStudy (mertsj):

Or is the problem the one you have posted which is: \[\frac{4}{x}-2=x-\frac{1}{x}-2\]

OpenStudy (mathmale):

If other users have to ask you for clarification regarding what you meant here, that's a sure sign that you should consider using parentheses to indicate which math operations are to be done first. How else is one to know whether you are dividing by x or by x-2?

jhonyy9 (jhonyy9):

my opinion is correct wrote how is above wrote by Mertsj so in this way the first step this recstriction of x not equal by 2 bc. than result denominator equal zero and hence the fraction is undefined - and the other way bc. denominators are same x-2 so just made equal the numerators and solve them for x so will result 4=x-1 x=5 hope this will help you